Circular No. 7906

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002db IN NGC 5683 B. Beutler and W. D. Li, University of California at Berkeley, for

the LOTOSS team (whose core members are Li, A. V. Filippenko, R. R. Treffers, and M. Schwartz, as of 2001 Sept. 1; see also IAUC 7514), report the discovery of an apparent supernova on an unfiltered CCD image taken with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) on May 18.3 UT (mag ∼ 17.6). The new object was confirmed on an image taken with the 0.6-m telescope at Beijing Astronomical Observatory by Y. L. Qiu and W. Zhou under poor conditions on May 20.6. SN 2002db is located at α = 14^{h}34^{m}51^{s}.63, δ = +48^{o}39′52′′.9 (equinox 2000.0), which is 8′′.0 west and 9′′.3 north of the nucleus of NGC 5683. A KAIT image taken on May 10.3 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2000fv IN MCG –04-35-11 Li also reports the discovery of an apparent supernova on unfiltered

CCD archival KAIT images taken during the course of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126). SN 2000fv is located at α = 14^{h}51^{m}09^{s}.98, δ = −26^{o}38′11′′.5 (equinox 2000.0), which is 4′′.2 east and 5′′.4 north of the nucleus of MCG –04-35-11. Unfiltered KAIT magnitudes for SN 2000fv: 2000 June 6.2 UT, 15.8; 24.2, 15.9; July 1.2, 16.0; 2002 May 14.4, [19.0:.

SAX J1819.3–2525 C. B. Markwardt, University of Maryland and Goddard Space Flight

Center (GSFC); and J. H. Swank, GSFC, report that RXTE PCA monitoring observations of the Galactic-center region on May 17.6 and 20.9 UT have shown that SAX J1819.3–2525 (IAUC 7119) = XTE J1819–254 (IAUC 7120) = V4641 Sgr (IAUC 7277) is again active as an x-ray source. This source was previously active in 1999 as a “microquasar” and is a confirmed dynamical blackhole system (Orosz et al. 2001, Ap.J. 555, 489). X-ray fluxes of 4.5 ± 0.3 and 10.0 ± 0.5 mCrab (in the band 2–10 keV) were detected at the two observation epochs, at a position consistent with the source. Follow-up RXTE observations are being scheduled.

SUPERNOVA 2002ct Corrigendum. On IAUC 7898, line 5, for 13′′ west read 13′′ east

2002 May 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green